OpenStudy (anonymous):

ok i'm confused right now......help! 1) Does a flat piece id aluminum foil sink or float? and why? 2) When squeezed into a ball, does it float or sink? why? Thanks !!

OpenStudy (anonymous):

We need to look at density and buoyancy forces. We know that an object floats in water when it is less dense than water. Aluminum has a density of \(2.70~ g / cm^3\) Water has a density of \(1.00~g/cm^3\) This means that a block of aluminum will sink because its density is greater than the density of water. What about a piece of aluminum foil? Let's take a piece as being 30cm x 30cm x 0.001 cm This sheet of aluminum foil would have a volume of \(0.9~ cm^3\). This sheet of aluminum foil will have a mass of \[m = 0.9*2.70 = 2.43g\] The mass of the water the aluminum foil displaces is \[m_w= 0.9 * 1.00 = 0.9 g\] The mass of the displaced water is less than the mass of the aluminum, therefore the sheet sinks. If we take the same sheet and crumple it up into a ball 5cm in diameter, the same 2.43 grams of aluminum now occupies 19.63 \(cm^3\) of volume. If we displace 19.63 \(cm^3\) of water, we have a mass of 19.63 grams. This is much more than the mass of the aluminum ball, therefore the ball floats. This same principle explains how giant tankers can be made to float even when they are made out of steel, which is denser than water.

OpenStudy (anonymous):

Oops. I calculated the volume of the ball wrong. It will be 130.9 cm^3, but the principle still holds.

OpenStudy (anonymous):

thank you so much! :)